in a market for a given commodity the quantity demand was 14 units hen the market price was 6 birr, when price was increases to 8 birr the quantity demand was decreased to 12 units. what is the elasticity of demand ?


1
Expert's answer
2021-12-12T17:41:52-0500

Elasticity of demand is computed as

"E_d=\dfrac{Q_2-Q_1}{P_2-P_1}\times \dfrac{P_1+P_2}{Q_1+Q_2}\\[0.4cm]"

Therefore

"E_d=\dfrac{12-14}{8-6}\times \dfrac{8+6}{12+14}\\[0.4cm]
E_d=-1\times \dfrac{7}{13}\\[0.3cm]
E_d={-0.54}"